Paradise Tours, Inc. just paid a dividend of $5.75. Analysts expect the company's dividend to grow by 35% this year, by 20% in year 2, and at a constant rate of 6% in Year 3 and thereafter. The required rate of return on PTI's stock is 14.00%. What is the best estimate of the stock's current intrinsic value? Note: Enter your answer rounded off to two decimal points. Do not enter $ or comma in the answer box. For example, if your answer is $123.456 then enter as 123.46 in the answer box.
Answer #1

D1 = D0 x (1 + g1) = 5.75 x (1 + 35%) = 7.7625

D2 = D1 x (1 + g2) = 7.7625 x (1 + 20%) = 9.315

D3 = D2 x (1 + g) = 9.315 x (1 + 6%) = 9.8739

Horizon value of dividends from year 3 onward, at the end of year 2 = D(HV) = D3 / (Ke - g) = 9.8739 / (14% - 6%) = 123.42375

Hence, the best estimate of the stock's current intrinsic value = PV of all the future dividends

= D1 / (1 + Ke) + [D2 + D(HV)] / (1 + Ke)2 = 7.7625 / (1 + 14%) + (9.315 + 123.42375) / (1 + 14%)2 = $  108.95